Please help.
Subtract
4 1/3-1 1/4

Answer by checkley71(8403)

1 1/3-1 1/4 HERE WE NEED TO CHANCE THESE MIXED FRACTION TO IMPROPER FRACTIONS THUS:
4/3-5/4 NOW WE NEED TO FIND A COMMON DENOMINATOR. 3*4=12 WILL DO.
4/3=(4*4)/12=16/12)
5/4=(5*3)/12=15/12 NOW WE CAN SUBTRACT THEM.
(16-15)/12=1/12 IS THE ANSWER.